OpenStudy (shamil98):

How to write the equation of the parabola, in vertex form, with focus at (2,-4) and directrix y = -6?

OpenStudy (anonymous):

Do uk the vertex form equation?

OpenStudy (shamil98):

y = a(x-h)^2 + k ?

OpenStudy (anonymous):

yes

OpenStudy (shamil98):

uh, what next? i thought the vertex was (h,k), not the focus.

OpenStudy (anonymous):

yea one sec

OpenStudy (anonymous):

give me like 5mins brb

OpenStudy (nincompoop):

laughing out loud

OpenStudy (shamil98):

HE DITCHED MEEEE

OpenStudy (nincompoop):

remember the relationship of the conics form and vertex form I showed you?

OpenStudy (shamil98):

4p = 1/a ?

OpenStudy (nincompoop):

the whole equations

OpenStudy (nincompoop):

where is the focus and where is the vertex

OpenStudy (shamil98):

4p(y – k) = (x – h)^2

OpenStudy (nincompoop):

p = 1/4 is the distance of the directrix on the axis of symmetry

OpenStudy (nincompoop):

directrix to the vertex

OpenStudy (shamil98):

explain it to me like im an 8th grader pls, im dumb

OpenStudy (nincompoop):

laughing my arse off you're not

OpenStudy (shamil98):

p = 1/4 is the distance of directrix on the axis of symmetry so.. what do i do with all this crap

OpenStudy (nincompoop):

I missed to write distance of vertex to directrix

OpenStudy (nincompoop):

let me ask you first, can you identify where does the parabola opens?

OpenStudy (shamil98):

axis of symmetry , y = -4 ?

OpenStudy (shamil98):

\[\huge \sqrt{(x_0 - 2)^2 + (y_0 +4)^2} = (y_0 - 2)\] \[\huge (x_0 - 2)^2 + (y_0 +4)^2 = (y_0 - 2)^2\] \[\huge x^2-4x+4 + y^2+8y+16 = y^2 +12y + 36\] \[\huge y = \frac{ x^2 }{ 4 } -x - 4\]

OpenStudy (shamil98):

ugh that should (y_0+6**

OpenStudy (shamil98):

y + 4 = x^2/4 - x (y+4) = x^2/4 - x but it isn't one of the answer choices -.-

OpenStudy (shamil98):

i could simplify it a bit more..

OpenStudy (shamil98):

y+4 = 1/4(x-2)^2 - 1 so 4(y+5) = (x-2)^2 which is actually an answer choice, so yeah i got it right ;D

OpenStudy (shamil98):

yo nin u dere